Akira Imada

Anadolu University, Electrical and Electronics Engineering Department, Iki Eylul Campus, Eskisehir 26470 Turkey

March 28, 2006

GECCO-2000, c/o AAAI, 445 Burgess Drive, Menlo Park, CA 94025, USA. Phone 650-328-3123.

Dear Sir,

Enclosed are eight copies of our paper submitted to the GECCO’2000 conference:

• title:

A Study of Hill Shape by Reverse EP Hill Climbing:

Why does Hebbian Peak Resist to be Found?

• author:

Akira Imada

Thank you.

With best regards,

Akira Imada